OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2) My answer is y - 7 = -3(x + 2) Am I right? :)

jimthompson5910 (jim_thompson5910):

you mixed up x and y

OpenStudy (johnweldon1993):

y - y1 = m(x - x1) remember? (7 , -2) (x y) try it again...

OpenStudy (anonymous):

y - 2 = -3(x + 7)

OpenStudy (johnweldon1993):

Remember your 'y' is a negative number...so if you have y - y1 = m(x - x1) y - (-2)...this becomes...?

OpenStudy (anonymous):

Oh yea!! Thats right, so it would be y + 2 = -3(x - 7)

OpenStudy (johnweldon1993):

There we go...good job....just try to remember these formulas you have been using...

OpenStudy (anonymous):

thank you! Sorry about that I'm getting tired.

OpenStudy (anonymous):

I'm going to write them down and study them!